ECON 402 HW 4 Page 3 5 Durable Goods (30 points) Tony is a seller of goods that last at most two periods. There are three potI need help with this problem
Answer #1

Three potential customers, H, L1, and L2, each of whom wants at most one good that is if good is purchased by a customer in the first period then she will not buy it in the subsequent period. Tony will charge price such that his payoff is maximized.

a) Subgame perfect Nash equilibrium is a refinement of Nash equilibrium used in dynamic games. Backward induction is used in determining subgame perfect equilibria in the case of finite games. First, consider the last actions of the game and find out which actions the final mover will take in order to maximize her utility. Here, the final set of actions is the set of price levels in period 2. And after that, the same process will continue for the second last set of actions. Here, the second last set of actions is all possible set of prices that can be charged in period 1.

Assume, if indifferent between buying and not buying, the customer will buy.

In period 2, price (p2) will be such that customer L1 and L2 purchase the good. So, the price will be 200 and payoff to the seller would be 200*2

In the first period, price charged will be such that customer H will purchase the good. Price (p1) would be 1400 and H will be indifferent between purchasing it in the first period and the second period. so, she will buy it in the first period and payoff to the seller would be 1400*1

(p1,p2) = (1400,200) is SPNE as this is calculated by the method of backward induction.

Tony's payoff = 1400*1 + 200*2 = 1800

Tony's payoff in period 1 = 1800

b) Tony has only two goods to sell at the beginning of the first period. It implies that he can sell it to only two customers. So, Tony would like to sell it to H and either L1 or L2 to maximize her payoff. So, the prices charged in two periods will remain the same that is (p1,p2) = (1400,200) but Tony's payoff will fall to 1400*1 + 200*1 = 1600

Tony's payoff in period 2= 1600